8-19 FIXED MANUFACTURING OVERHEAD VARIANCE ANALYSIS. The Lebanese Bakery Inc. also allocates FMOH to products on...

  1. 8-19 FIXED MANUFACTURING OVERHEAD VARIANCE ANALYSIS. The Lebanese Bakery Inc. also allocates FMOH to products on the basis of standard direct manufacturing labour-hours. For 2018, FMOH was budgeted at $4.00 per DMLH. Actual FMOH incurred during the year was $272,000.

    LO 1

    1. Rate variance, $35,200 F

    Baguettes are baked in batches of 100 loaves. Following are some pertinent data for Lebanese Bakery Inc.:

    Direct manufacturing labour use 2.00 DMLH per batch
    Fixed manufacturing overhead $4.00 per DMLH

    Lebanese Bakery Inc. recorded the following additional data for the year ended December 31,2018:

    Planned (budgeted) output 3,840,000 baguettes
    Actual production 3,360,000 baguettes
    Direct manufacturing labour 50,400 DMLH
    Actual fixed MOH $272,000

Required

  1. Prepare a variance analysis of FMOH costs.

  2. Is fixed overhead under- or overallocated? By how much?

  3. Comment on your results. Discuss the various variances, and explain what may be driving them
Answer #1

Part 1

Budgeted standard direct manufacturing labor used = 2/100 =0.02 per baguette

Budgeted output = 3,840,000 baguettes

Budgeted standard direct manufacturing labor-hours = 3,840,000 × 0.02= 76800 hours

Budgeted fixed manufacturing overhead costs = 76800 × $4.00 per hour= $307200

Actual output = 3360000 baguettes

Allocated fixed manufacturing overhead = 3360000 × 0.02 × $4= $268800

Actual Costs Incurred (1)

Same Budgeted Lump Sum (as in Static Budget) Regardless of Output Level (2)

Flexible Budget: Same Budgeted Lump Sum (as in Static Budget) Regardless of Output Level (3)

Allocated: Budgeted Input Qty. Allowed for Actual Output × Budgeted Rate(4)

$272000

$307200

$307200

$268800

$35200 F

Spending variance

Never a variance

$38400 U

Production-volume

$35200 F

Flexible-budget variance

$38400 U

Production-volume

$3200 U

Underallocated fixed overhead

(Total fixed overhead variance)

Part 2

Fixed manufacturing overhead is underallocated by $3200.

Part 3

The unfavorable production-volume variance of $38400 indicates the difference between the budgeted 3,840,0000 baguettes and the lower actual 3,360,000 baguettes produced—the fixed cost capacity not used. The favorable spending variance of $35200 indicates that the actual aggregate of fixed costs of $272,000 is lower than the budget amount $307200.